OpenStudy (anonymous):

Simplify (2sqrt5 + 3sqrt7)^2?? I keep getting different answers every-time I try to solve. What steps do I do??

OpenStudy (zepp):

Show what you did?

OpenStudy (zepp):

Okay, since 2sqrt(5) and 3sqrt(7) are annoying, can you develop them so it would be only sqrt(#)?

OpenStudy (zepp):

E.g.: 2sqrt(5) could be expanded to sqrt(20)

OpenStudy (anonymous):

3sqrt(7) = 42?

OpenStudy (zepp):

Nope, 3 = sqrt(9) sqrt(9)*sqrt(7) = sqrt(63)

OpenStudy (zepp):

\[(\sqrt{20}+\sqrt{63})^2\]

OpenStudy (zepp):

Do something with it Hint: \[(a+b)^2 = a^2+2ab+b^2\]

OpenStudy (anonymous):

So sqrt(20) is A?

OpenStudy (zepp):

Yes.

OpenStudy (anonymous):

{ sqrt(20) + sqrt(63)}^2 = sqrt(20)^2 + 2(20)(63) + (63)^2

OpenStudy (zepp):

Uh, you forgot some sqrt somewhere

OpenStudy (anonymous):

{( 2sqrt(5) + 3sqrt(7)^2} = 83 + 12sqrt(35)

OpenStudy (zepp):

\[(\sqrt{20})^2+2(\sqrt{20}\sqrt{63})+(\sqrt{63})^2\]

OpenStudy (zepp):

So it would be 20 and 63 at both ends, since sqrt and ^2 cancels out \[20 + 2\sqrt{20*63}+63\] Answer should be pretty obvious now, simplify this and you'll get the answer.

OpenStudy (anonymous):

83 + 12 sqrt(35)

OpenStudy (zepp):

wooot! :D